When two identical ions are separated by a distance of 2.0×10−10 m , the electrostatic force each exerts on the other is 5.6×10−9 N .

Part A

How many electrons are missing from each ion?

Express your answer as an integer.

Ne =
Answer #1

Here ,

let the number of electrons missing is n

charge on each , q = n * e = n * 1.602 *10^-19

Now , for the force

5.6 *10^-9 = 9 *10^9 * (n * 1.602 *10^-19)^2/(2 *10^-10)^2

solving for n

n = 1

the number of electons missing from each is 1